40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16748     Entry time: Tue Mar 29 17:35:54 2022
Author: Paco 
Type: Update 
Category: SUS 
Subject: Damping fix on BS, AS4, PR2, and PR3 

[Ian, Paco]

  • We removed the "cheby" filters from AS4, PR2 and PR3 which had been misplaced after copying from the old SUS models. After removing them, the new SOS damped fine. Note that because of the Input matrices, the filters have to be enabled all at once for the MIMO loop to make sense.
  • We also disabled the "Cheby" filter on BS and saw it damp better. We don't understand this yet, but perhaps it's just a consequence of the many changes in the BSC that have rendered this filter obsolete.
  • we also reduced the damping gains on PR2, PR3 and AS4 to prevent overflow values. After the adjustments the optics were damping fine.
ELOG V3.1.3-